Use inductive reasoning to predict the most probable next number in each of the following lists.


A.5, 10, 15, 20, 25, ? (?)


1
Expert's answer
2022-05-13T02:24:03-0400

10-5=5

15-10=5

20-15=5

25-20=5

Common difference is 5.

So the next number is 30 because 30-25=5.